OpenStudy (anonymous):

I have a circuit with the following: R1=1.0kohms R2= 3.0 kohms R3= 1.25kohms r4=2.2kohms R5=1.8kohm R6=970 OHMS R7=1.7kohms R8=5kohms with keyA=20% Voltage 1=12V and Voltage 2=3 Volts. I am asked to find the KVL equation.

OpenStudy (anonymous):

we need the circuit mate!

OpenStudy (anonymous):

I'll provide a simple example that should help you get on the way with finding KVL equations.  What Kirchoff said was that the sum of all voltages in a loop must be zero. That means that `-V1 + Vr1 + Vr2 = 0` and `-V2 + Vr2 + Vr3 = 0`. Now the problem is the values of Vr1, Vr2 and Vr2. Using Ohm's law, you know that `V=IR`, so `Vr1 = Ir1 * R1`, `Vr2 = Ir2 * R2` and `Vr3 = Ir3 * R3`. Also, note that `Ir2 = Ir1 + Ir3` (Sorry, didn't include those currents in my drawing. Ir1 flows through R1 from left to right and Ir3 flows through R3 from right to left). Now we know that `Vr2 = Ir2 * R2 = ( Ir1 + Ir3 ) * R1` and we can add that information to our first two equations: \[V_1 = V_{R_1} + V_{R_2} = I_{R_1} \cdot R_1 + (I_{R_1} + I_{R_3})\cdot R_2\] \[V_2 = V_{R_2} + V_{R_3} = (I_{R_1} + I_{R_3}) \cdot R_2 + I_{R_3} \cdot R_3\] With some rewriting we get the KVL equations (which we can solve for the two currents): \[V_1 = (R_1 + R_2) \cdot I_{R_1} + R_2 \cdot I_{R_3}\] \[V_2 = R_2 \cdot I_{R_1} + (R_2 + R_3) \cdot I_{R_3}\] Now try to apply this on your own circuit. I hope this helps.